OpenStudy (anonymous):

Can somebody help me find the value for 'X' in this equation? 10^x-7 = 10^3x+8

OpenStudy (johnweldon1993):

Since these "exponents" both have the same base...we can ignore that 10....and just set the exponents equal to each other \[\large x - 7 = 3x + 8\] Because when we find out what 'x' equals we can make both 10^(that) and they will also be equal

OpenStudy (johnweldon1993):

Assuming this is \[\large 10^{x - 7} = 10^{3x + 8}\]?

OpenStudy (anonymous):

Yes sir.

OpenStudy (johnweldon1993):

Alright...so yes all we need to do is solve that top equation \[\large x - 7 = 3x + 8\] subtract 8 from both sides and subtract x from both sides \[\large 2x = -15\] \[\large x = ?\]

OpenStudy (anonymous):

X would be -7.5 is that correct?

OpenStudy (johnweldon1993):

Correct...now lets see if it IS correct \[\large 10^{(-7.5 - 7)} = 10^{3(-7.5) + 8}\] \[\large 10^{-14.5} = 10^{-14.5}\] \[\large \color \red{\checkmark}\]

OpenStudy (anonymous):

Thank you so much for the help.

OpenStudy (johnweldon1993):

No problem!